# Riemannian Zeroth-Order Gradient Estimation with Structure-Preserving Metrics for Geodesically Incomplete Manifolds

### v1

@@ 1 Introduction
where $(\mathcal{M},g)$ forms a $d$-dimensional Riemannian manifold, the individual loss $f(\cdot;\xi):\mathcal{M}\to\mathbb{R}$ is a smooth function depending on a random data point $\xi$ drawn from a distribution $\Xi$. The Riemannian metric $g$ allows us for defining the first-order gradient $\nabla f(p;\xi)$ in the tangent space at each $p\in\mathcal{M}$, leading to the standard first-order Riemannian stochastic gradient method (Ring and Wirth, 2012; Bonnabel, 2013; Smith, 2014; Sato, 2021).
In many practical scenarios, especially when the system incorporates non-differentiable external solvers or black-box objective functions especially when dealing with non-differentiable modules or black-box objective functions, the explicit gradient of the objective function is either unavailable or prohibitively expensive to compute. This practical challenge necessitates the use of zeroth-order optimization technique to approximate the gradient direction solely using the function evaluation (Nesterov and Spokoiny, 2017; Li et al., 2023a), given by
@@ 1.1 Challenges in Riemannian Zeroth-Order Optimization
While existing analyses of Riemannian zeroth-order optimization establish convergence guarantees under various algorithms and assumptions (Chattopadhyay et al., 2015; Fong and Tiňo, 2019; Wang et al., 2021a; Wang and Feng, 2022; Maass et al., 2022; Nguyen and Balasubramanian, 2023; Li et al., 2023a; b; Wang, 2023; He et al., 2024; Wang et al., 2023; Goyens et al., 2024; Zhou et al., 2025; Ochoa and Poveda, 2025), a fundamental yet often overlooked issue arises from the local nature of the exponential map (or, more generally, retractions).
In practice, Riemannian zeroth-order methods often endow $\mathcal{M}$ with an Euclidean metric $g_{E}$ by viewing it as a submanifold of an ambient Euclidean space $\mathbb{R}^{n}$ and inheriting the metric from the embedding. This setting helps simplify numerical computations, but it has a fundamental limitation: the inherited Euclidean metric $g_{E}$ may not be geodesically complete. Specifically, for a point $p\in\mathcal{M}$, the exponential map $\exp_{p}$ is not necessary globally defined over the entire tangent space $T_{p}\mathcal{M}$. Consequently, a randomly sampled tangent vector $v\in T_{p}\mathcal{M}$ may fall outside the domain of $\exp_{p}$, making $\exp_{p}(v)$ undefined.
Theoretically, one could instead begin with a geodesically complete metric, under which the exponential map $\exp:T\mathcal{M}\to\mathcal{M}$ is globally defined on the full tangent bundle $T\mathcal{M}$. The Nomizu-Ozeki theorem (Nomizu and Ozeki, 1961; Lee, 2018) guarantees the existence of such a complete metric on any smooth manifold without boundary. Then by applying the Nash embedding theorem (Nash, 1956), one could, in principle, obtain an equivalent geodesically complete Euclidean metric, allowing direct application of existing convergence analyses. However, the constructive proof of Nash’s theorem is numerically nontrivial, making it infeasible for practical optimization algorithms.
@@ 1.2 Contributions
Contribution 1 (Structure-Preserving Metric Construction): To address the potential geodesic incompleteness of the given metric $g$, we construct the structure-preserving metrics $g^{\prime}$ (Definition˜2.5) in Theorem˜2.6 that: (i) is geodesically complete, (ii) is conformally equivalent to the original metric $g$, and (iii) ensures any $\epsilon$-stationary point under $g$ is also an $\epsilon$-stationary point under $g^{\prime}$. These properties allow us to work with the new metric $g^{\prime}$ while maintaining the desired property as the original metric $g$.
However, adopting the structure-preserving metric raises a fundamental challenge: the geometry induced by $g^{\prime}$ generally differs from that of $g$. In particular, $g^{\prime}$ is typically no longer an Euclidean metric inherited from the original ambient Euclidean space, which precludes the direct use of standard Riemannian zeroth-order gradient estimators (Li et al., 2023b; a). Overcoming this mismatch between estimator design and underlying geometry leads to our second contribution.
@@
1.3 Applications of Riemannian zeroth-order optimization on geodesically incomplete manifolds
@@ 2 Main Results
2.1 Assumptions
In this subsection, we summarize the assumptions used in our theoretical analysis. A brief manifold preliminary is included in Appendix˜B. Detailed discussions of each assumption are provided in Section˜C.1.
@@ 2.1 Assumptions
$f(\cdot;\xi):\mathcal{M}\to\mathbb{R}$
satisfies the following two properties:
  1. (a)
    $L$-Bounded Hessian; for all $p\in\mathcal{M}$,
  2. (b)
@@
There exist constants $\rho>0$ and $M_{3},M_{4}>0$ such that
@@ 2.2 Structure-Preserving Metric
Here, we include a brief discussion on the motivation for introducing each condition.
  • The first condition (geodesic completeness) ensures that if we set the perturbation stepsize $\mu<\rho$ and fix the random vector $v$ on the $g$-unit sphere $\mathbb{S}^{d-1}\subset T_{p}\mathcal{M}$, the perturbed point $\mu v\in T_{p}\mathcal{M}$ will always be within the domain of the exponential map.
  • The conformal equivalence condition preserves the set of stationary points; that is, for any smooth function $f:\mathcal{M}\to\mathbb{R}$, if $p$ is a stationary point under $g$, then it is also a stationary point under $g^{\prime}$, and vice versa.
  • The $\epsilon$-stationarity preservation condition gives rise to the name “stationary-preserving metric”. It states that any $\epsilon$-stationary point under $g$ remains an $\epsilon$-stationary point under $g^{\prime}$, ensuring that the transformation leaves the original set of $\epsilon$-stationary points unchanged. We emphasize, however, that the converse need not hold: an $\epsilon$-stationary point under $g^{\prime}$ is generally not an $\epsilon$-stationary point under $g$. Nevertheless, under suitable conditions, this asymmetry does not affect the overall complexity guarantees as we will discuss it in Corollary˜2.10.
@@ 2.3 Intrinsic Zeroth-Order Gradient Estimation under Non-Euclidean Metric
Let $(\mathcal{M},g)$ be a complete $d$-dimensional Riemannian manifold and $p\in\mathcal{M}$. Let $f:\mathcal{M}\to\mathbb{R}$ be a smooth function and suppose that Assumptions˜2.3 and 2.4 hold. Fix a perturbation stepsize $\mu>0$ satisfying $\mu^{2}\leq\min\{\frac{1}{d-1},\frac{1}{2}+\frac{6}{d}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{8}{d^{2}}}\},$ and for any unit vector $v\in T_{p}\mathcal{M}$ define the symmetric zeroth-order estimator as in Equation˜3. Then, for $v\sim\mathrm{Unif}(\mathbb{S}^{d-1})$ uniformly sampled from the $g_{p}$-unit sphere in $T_{p}\mathcal{M}$,
@@ 2.5 Convergence of Zeroth-Order SGD under Non-Euclidean Metric
Let $(\mathcal{M},g)$ be a geodesically complete $d$-dimensional Riemannian manifold. Let $f:\mathcal{M}\to\mathbb{R}$ be a smooth function and suppose that Assumptions˜2.1, 2.2, 2.3 and 2.4 hold. Define the symmetric zeroth-order estimator as in Equation˜5. Let $\{p_{t}\}_{t=1}^{T}$ be the SGD dynamic finding the stationary point of Equation˜1 generated by the update rule Equation˜4 with requiring $\eta\lesssim\sqrt{\frac{d}{T}}$ and $\mu^{2}\lesssim\sqrt{\frac{d}{T}}$ (explicitly specified in Equation˜21), then there exists constants $\mathsf{C}_{1},\mathsf{C}_{2},\mathsf{C}_{3}>0$ such that
@@ 3 Experiments
In the experimental section, we aim to validate the theoretical findings presented in Section˜2. The two synthetic experiments are designed to investigate the following questions:
  1. (i)
    How does sampling bias influence the convergence behavior of Riemannian zeroth-order SGD?
  2. (ii)
    How does the curvature of the underlying manifold affect the accuracy of gradient estimation?
In addition, we conduct a real-world experiment on mesh optimization (Hoppe et al., 1993; Belbute-Peres et al., 2020; Ma et al., 2025), a practical application in which the positions of nodes are naturally represented as points on the probability simplex. All source codes are attached to the supplementary.
@@ 3.3 Gradient-Based Mesh Optimization
where $\nabla^{2}$ denotes the Laplace operator, $k=10$ is the wave number, and $f$ is the eigenfunction. The ground-truth solution is computed on a fine mesh with resolution $200\times 200$. Our goal is to optimize the node positions of a regular coarse mesh with resolution $20\times 20$ so that its performance approximates that of the ground-truth solution.
The mesh node (in our setting, boundary nodes are fixed and excluded from optimization) is represented using a simplex formulation: each trainable node $p=(x,y)$ is expressed as a convex combination of its six neighbors under the regular triangular initialization. This parameterization naturally leads to a manifold optimization problem. However, the coordinate simplex, under its canonical embedding, is geodesically incomplete. To ensure the exponential map remains well-defined and to prevent perturbed nodes from crossing mesh edges, we adopt our proposed structure-preserving approach and compare it against several natural baselines, as illustrated in Figure˜5.
@@ C.1 Assumptions
    $L$-Bounded Hessian; for all $p\in\mathcal{M}$,
@@ C.2 Supporting Lemmas
Let $\mathcal{M}$ be a smooth manifold. Then there exists a smooth function $\rho:\mathcal{M}\to[0,+\infty)$ that is proper; that is, for every compact set $C\subset\mathbb{R}$, $\rho^{-1}(C)$ is compact in $\mathcal{M}$.
@@
    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\overset{(i)}{\leq}}\frac{\mu^{4}M_{3}^{2}}{9d^{2}(d+2)^{2}}+\frac{\mu^{6}}{144}M^{2}_{4}+\frac{\mu^{4}}{(d+2)^{2}}\kappa^{2}\|\nabla f(p_{t};\xi_{t})\|^{2}_{p_{t}}$
@@
$\mu^{2}\leq\min\{\frac{1}{d-1},\frac{1}{2}+\frac{6}{d}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{8}{d^{2}}}\},$

### v2

@@ 1 Introduction
where $(\mathcal{M},g)$ forms a $d$-dimensional Riemannian manifold, the individual loss $f(\cdot;\xi):\mathcal{M}\to\mathbb{R}$ is a smooth function depending on a random data point $\xi$ drawn from a distribution $\Xi$. The Riemannian metric $g$ allows us for defining the first-order gradient $\nabla f(p;\xi)$ in the tangent space at each $p\in\mathcal{M}$, leading to the standard first-order Riemannian stochastic gradient method (Ring and Wirth, 2012; Bonnabel, 2013; Smith, 2014; Sato, 2021). In many practical scenarios, especially when the system incorporates non-differentiable external solvers or black-box objective functions especially when dealing with non-differentiable modules or black-box objective functions, the explicit gradient of the objective function is either unavailable or prohibitively expensive to compute. This practical challenge necessitates the use of zeroth-order optimization technique to approximate the gradient direction solely using the function evaluation (Nesterov and Spokoiny, 2017; Li et al., 2023a), given by
@@ 1.1 Challenges in Riemannian Zeroth-Order Optimization
While existing analyses of Riemannian zeroth-order optimization establish convergence guarantees under various algorithms and assumptions (Chattopadhyay et al., 2015; Fong and Tiňo, 2019; Wang et al., 2021a; Wang and Feng, 2022; Maass et al., 2022; Nguyen and Balasubramanian, 2023; Li et al., 2023a; b; Wang, 2023; He et al., 2024; Wang et al., 2023; Goyens et al., 2024; Zhou et al., 2025; Ochoa and Poveda, 2025), a fundamental yet often overlooked issue arises from the local nature of the exponential map (or, more generally, retractions). In practice, Riemannian zeroth-order methods often endow $\mathcal{M}$ with an Euclidean metric $g_{E}$ by viewing it as a submanifold of an ambient Euclidean space $\mathbb{R}^{n}$ and inheriting the metric from the embedding. This setting helps simplify numerical computations, but it has a fundamental limitation: the inherited Euclidean metric $g_{E}$ may not be geodesically complete. Specifically, for a point $p\in\mathcal{M}$, the exponential map $\exp_{p}$ is not necessary globally defined over the entire tangent space $T_{p}\mathcal{M}$. Consequently, a randomly sampled tangent vector $v\in T_{p}\mathcal{M}$ may fall outside the domain of $\exp_{p}$, making $\exp_{p}(v)$ undefined. Theoretically, one could instead begin with a geodesically complete metric, under which the exponential map $\exp:T\mathcal{M}\to\mathcal{M}$ is globally defined on the full tangent bundle $T\mathcal{M}$. The Nomizu-Ozeki theorem (Nomizu and Ozeki, 1961; Lee, 2018) guarantees the existence of such a complete metric on any smooth manifold without boundary. Then by applying the Nash embedding theorem (Nash, 1956), one could, in principle, obtain an equivalent geodesically complete Euclidean metric, allowing direct application of existing convergence analyses. However, the constructive proof of Nash’s theorem is numerically nontrivial, making it infeasible for practical optimization algorithms.
@@ 1.2 Contributions
Contribution 1 (Structure-Preserving Metric Construction): To address the potential geodesic incompleteness of the given metric $g$, we construct the structure-preserving metrics $g^{\prime}$ (Definition˜2.5) in Theorem˜2.6 that: (i) is geodesically complete, (ii) is conformally equivalent to the original metric $g$, and (iii) ensures any $\epsilon$-stationary point under $g$ is also an $\epsilon$-stationary point under $g^{\prime}$. These properties allow us to work with the new metric $g^{\prime}$ while maintaining the desired property as the original metric $g$. However, adopting the structure-preserving metric raises a fundamental challenge: the geometry induced by $g^{\prime}$ generally differs from that of $g$. In particular, $g^{\prime}$ is typically no longer an Euclidean metric inherited from the original ambient Euclidean space, which precludes the direct use of standard Riemannian zeroth-order gradient estimators (Li et al., 2023b; a). Overcoming this mismatch between estimator design and underlying geometry leads to our second contribution.
@@
1.3 Applications
@@ 2 Main Results
Here we summarize the assumptions used in our theoretical analysis. A brief manifold preliminary is included in Appendix˜B. Detailed discussions of each assumption are provided in Section˜C.1.
@@
In the optimization problem given by Equation˜1, the individual loss function $f(\cdot;\xi):\mathcal{M}\to\mathbb{R}$ satisfies the following two properties: (a) $L$-Bounded Hessian; for all $p\in\mathcal{M}$, the Hessian matrix at the point $p$ is bounded by $L$. (b) Lower boundedness; the infimum $f_{\xi}^{*}:=$ exists almost surely with $\xi\sim\Xi$.
@@
There exist constants $\rho>0$ and $M_{3},M_{4}>0$ such that $\bigl\|\nabla^{3}f(q)\bigr\|_{\mathrm{HS}}\leq M_{3}$ and $\bigl\|\nabla^{4}f(q)\bigr\|_{\mathrm{HS}}\leq M_{4}$ for all $q\in\mathcal{B}_{p}(p,\rho)$, where $\mathcal{B}_{p}(p,\rho)$ denotes the geodesic ball of radius $\rho$ and $\|\cdot\|_{\mathrm{HS}}$ is the Hilbert-Schmidt norm.
@@
There exists a constant $\kappa\geq 0$ such that the sectional curvature of the Riemannian manifold $(\mathcal{M},g)$ satisfies $|K_{p}(\sigma)|\leq\kappa,$ for every point $p\in\mathcal{M}$ and every $2$-plane $\sigma\subset T_{p}\mathcal{M}$. Equivalently, $\kappa\leq K_{p}(\sigma)\leq\kappa$ for all $p$ and $\sigma$.
@@ 2.1 Structure-Preserving Metric
Here, we include a brief discussion on the motivation for introducing each condition: (i) The first condition (geodesic completeness) ensures that if we set the perturbation stepsize $\mu<\rho$ and fix the random vector $v$ on the $g$-unit sphere $\mathbb{S}^{d-1}\subset T_{p}\mathcal{M}$, the perturbed point $\mu v\in T_{p}\mathcal{M}$ will always be within the domain of the exponential map. (ii) The conformal equivalence condition preserves the set of stationary points; that is, for any smooth function $f:\mathcal{M}\to\mathbb{R}$, if $p$ is a stationary point under $g$, then it is also a stationary point under $g^{\prime}$, and vice versa. (iii) The $\epsilon$-stationarity preservation condition gives rise to the name “stationary-preserving metric”. It states that any $\epsilon$-stationary point under $g$ remains an $\epsilon$-stationary point under $g^{\prime}$, ensuring that the transformation leaves the original set of $\epsilon$-stationary points unchanged. We emphasize, however, that the converse need not hold: an $\epsilon$-stationary point under $g^{\prime}$ is generally not an $\epsilon$-stationary point under $g$. Nevertheless, under suitable conditions, this asymmetry does not affect the overall complexity guarantees as we will discuss it in Corollary˜2.10.
@@ 2.2 Intrinsic Zeroth-Order Gradient Estimation under Non-Euclidean Metric
Let $(\mathcal{M},g)$ be a complete $d$-dimensional Riemannian manifold and $p\in\mathcal{M}$. Let $f:\mathcal{M}\to\mathbb{R}$ be a smooth function and suppose that Assumptions˜2.3 and 2.4 hold. Fix a perturbation stepsize $\mu>0$ satisfying $\mu^{2}\leq\min\{\frac{1}{d-1},\frac{1}{2}+\frac{6}{d}+\frac{8}{d^{2}}\},$ and for any unit vector $v\in T_{p}\mathcal{M}$ define the symmetric zeroth-order estimator as in Equation˜3. Then, for $v\sim\mathrm{Unif}(\mathbb{S}^{d-1})$ uniformly sampled from the $g_{p}$-unit sphere in $T_{p}\mathcal{M}$,
@@ 2.4 Convergence of Zeroth-Order SGD under Non-Euclidean Metric
Let $(\mathcal{M},g)$ be a geodesically complete $d$-dimensional Riemannian manifold. Let $f:\mathcal{M}\to\mathbb{R}$ be a smooth function and suppose that Assumptions˜2.1, 2.2, 2.3 and 2.4 hold. Define the symmetric zeroth-order estimator as in Equation˜5. Let $\{p_{t}\}_{t=1}^{T}$ be the SGD dynamic finding the stationary point of Equation˜1 generated by the update rule Equation˜4 with requiring $\eta\lesssim\sqrt{\frac{d}{T}}$ and $\mu^{2}\lesssim\sqrt{\frac{d}{T}}$ (specified in Equation˜21), then there exists constants $\mathsf{C}_{1},\mathsf{C}_{2},\mathsf{C}_{3}>0$ such that
@@ 3 Experiments
In the experimental section, we aim to validate the theoretical findings presented in Section˜2. The two synthetic experiments are designed to investigate the following questions: (i) How does sampling bias influence the convergence behavior of Riemannian zeroth-order SGD? (ii) How does the curvature of the underlying manifold affect the accuracy of gradient estimation? In addition, we conduct a real-world experiment on mesh optimization (Hoppe et al., 1993; Belbute-Peres et al., 2020; Ma et al., 2025), a practical application in which the positions of nodes are naturally represented as points on the probability simplex.
@@ 3.3 Gradient-Based Mesh Optimization
where $\nabla^{2}$ denotes the Laplace operator, $k=10$ is the wave number, and $f$ is the eigenfunction. The ground-truth solution is computed on a fine mesh with resolution $200\times 200$. Our goal is to optimize the node positions of a regular coarse mesh with resolution $20\times 20$ so that its performance approximates that of the ground-truth solution. The mesh node (in our setting, boundary nodes are fixed and excluded from optimization) is represented using a simplex formulation: each trainable node $p=(x,y)$ is expressed as a convex combination of its six neighbors under the regular triangular initialization. This parameterization naturally leads to a manifold optimization problem. However, the coordinate simplex, under its canonical embedding, is geodesically incomplete. To ensure the exponential map remains well-defined and to prevent perturbed nodes from crossing mesh edges, we adopt our proposed structure-preserving approach and compare it against several natural baselines, as illustrated in Figure˜5.
@@ C.1 Assumptions
    $L$-Bounded Hessian; for all $p\in\mathcal{M}$, the Hessian matrix at the point $p$ is bounded by $L$.
@@ C.2 Supporting Lemmas
Let $\mathcal{M}$ be a smooth manifold. Then there exists a smooth function $\rho:\mathcal{M}\to[0,+\infty)$ that is proper; i.e., for every compact set $C\subset\mathbb{R}$, $\rho^{-1}(C)$ is compact in $\mathcal{M}$.
@@
    $\overset{(i)}{\leq}\frac{\mu^{4}M_{3}^{2}}{9d^{2}(d+2)^{2}}+\frac{\mu^{6}}{144}M^{2}_{4}+\frac{\mu^{4}}{(d+2)^{2}}\kappa^{2}\|\nabla f(p_{t};\xi_{t})\|^{2}_{p_{t}}$
@@
$\mu^{2}\leq\min\{\frac{1}{d-1},\frac{1}{2}+\frac{6}{d}+\frac{8}{d^{2}}\},$
@@ D.4 Additional Discussions on Stability
Appendix E Limitations and Future Work
Appendix F The Use of Large Language Models (LLMs)
In preparing this manuscript, we employed Large Language Models (LLMs) as general-purpose assistive tools in the following ways:
  • Literature review support. We used the Deep Research functionality provided by existing AI platforms to help gather references and draft preliminary summaries of related work.
  • Language refinement. We used AI chatbots hosted on multiple platforms to generate the abstract and to improve the clarity, style, and readability of the manuscript.
  • Proof verification. We used AI chatbots to check the logical consistency, correctness, and completeness of our formal proofs.
  • Codes Generation. We also applied the AI agent to generate a part of experimental codes.
All LLM-assisted outputs were critically reviewed, verified, and, where necessary, revised by the authors. We take full responsibility for the content of this manuscript. LLMs were not involved in generating research ideas, drawing scientific conclusions, or contributing original insights.